A car tire has a pressure of 30 psi at a temperature of 27ºC. Calculate the extremes of pressure  caused by temperatures ranging from -20.0ºC on a cold winter day to 50.0ºC while being driven  on a hot summer day.


1
Expert's answer
2021-02-01T03:51:57-0500

According to Gay-Lussac's Law, the pressure of a given mass of a gas is directly proportional to its absolute temperature, so long as volume is kept constant.

Thus,

P"\alpha"T

P = T x constant

P = Tk

"\dfrac{P}{T} = k"


Therefore,

"\dfrac{P1}{T1} = \dfrac{P2}{T2}"


Solving for P2 becomes

"P2 =\dfrac{P1T2}{T1}" T1=(27+273)K = 300K

P1 = 30psi


(i) During winter when T2 = (-20 + 273)K = 253K


P2 "=\dfrac{30psi x253K}{300K} = 25.3psi"


(ii) During winter when T2=(50 + 273)K = 323K


P2 "=\dfrac{30psi x 323K}{300K} = 32.3psi"